OpenStudy (anonymous):

A transformation T : (x, y) → (x + 3, y + 1). Find the preimage of the point (4, 3) under the given transformation.

OpenStudy (gorv):

so after transformation new point=x+3 and = y+1 so u need to find x and y

OpenStudy (gorv):

4=x+3 find x from here 3=y+1 find from here